OpenStudy (anonymous):

Need help with scientic notation!!! (3.4 x 10^3)(2.8 x 10^8)

OpenStudy (anonymous):

3.4 * 2.8 * 10^11 = 9.52 * 10^11

OpenStudy (anonymous):

Thank you!!!! :) I have two more. Can you help me by showing the work too? It would be so helpful. (5.8 x 10^-6)^4

OpenStudy (anonymous):

(5.8)^4 * (10^-6)^4 = (5.8)^4 * 10^(-6 * 4) = (5.8)^4 * 10^-24

OpenStudy (anonymous):

i have to go now, dont worry about my last one. Thanks for all ur help, ur awesome! :D